4. According to an IRS study, it takes a mean of 330 minutes for taxpayers to prepare, copy,

and electronically file a 1040 tax form. This distribution of times follows the normal distribution

and the standard deviation is 80 minutes. A consumer watchdog agency selects

a random sample of 40 taxpayers.

a. What is the standard error of the mean in this example?

b. What is the likelihood the sample mean is greater than 360 minutes?
Answer #1

(a)

The standard Error of the mean is given by:
SE = /

= 80/ = 12.6491

(b)

To find P( > 360):

Z = (360 - 330)/12.6491 = 2.37

Table of Area Under Standard Normal Curve gives area = 0.4911

So,

P( > 360) = 0.5 - 0.4911 = 0.0089